OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today?

OpenStudy (anonymous):

i should make a template of this i have seen it maybe 5 times a week for 2 years now

OpenStudy (anonymous):

Its from virtual school >.<

OpenStudy (johnweldon1993):

Well we can set up the 2 equations we have been given Let 1 step equations = 'o' Let 2 step equations = 't' \[o + t = 1070\] \[3o - 2t = 110\] We can solve that top equation for...lets do 't' \[t = 1070 - o\] And now we can substitute this into that second equation for 't' \[3o - 2(1070 - o) = 110\] and now simplify \[3o - 2140 + 2o = 110\] add 2140 to both sides \[3o - 2140 + 2140 + 2o = 110 +2140\] \[3o + 2o = 2250\] Combine like terms \[5o = 2250\] Divide both sides by 5 \[o = \frac{ 2250 }{ 5 }\] What does 'o' equal...?

OpenStudy (anonymous):

Way to confuse the blonde...

OpenStudy (johnweldon1993):

Lol "I'm blonde too :P" and is that confusing...??

OpenStudy (anonymous):

yes! hey i need help with another problem though...mind helpin?

OpenStudy (johnweldon1993):

Well, I think we should make sure you have a firm grasp on what happened here first...

OpenStudy (anonymous):

Oh lord...

OpenStudy (anonymous):

reminds me of a dylan album

OpenStudy (anonymous):

So confused...

OpenStudy (johnweldon1993):

Lol it's not bad I swear :) First lets highlight the important parts of the question "The number of one-step equations and two-step equations who have been eliminated today is equal to 1070! If three times the number of one-step equations minus twice the number of two-step equations is equal to 110, how many one-step equations auditioned today?" "The number of one-step equations and two-step equations who have been eliminated today is equal to 1070!" This means that the sum of both these variables is equal to 1070 right? Remember how I let one step equations be represented by 'o' and two step represented by 't'? so this can be written as o + t = 1070 "If three times the number of one-step equations minus twice the number of two-step equations is equal to 110" 3 times the one steps (3 x o = 3o) minus (-) twice the number of two step (2 x t = 2t) is 110 put altogether 3o - 2t = 110 How you doing so far?

OpenStudy (anonymous):

okay

OpenStudy (johnweldon1993):

Okay...this is called a system of equations...you have 2 equations to work with ...to solve your problem o + t = 1070 and 3o + 2t = 110 Now...you cannot possibly solve these equations "as is"....2 unknowns is too much! so what you do...is solve one of these equations...to isolate 1 variable This is where I showed that I would solve that top equation for 't' o + t = 1070 to solve for 't'...I want to get 't' by itself.....so it would be t = ....something so o + t = 1070 if I subtract an 'o' from both sides...I will only have 't' left on that left side...so I did that o + t = 1070 -o -o t = 1070 - o With me still...?

OpenStudy (anonymous):

yup

OpenStudy (johnweldon1993):

typo up there should be 3o - 2t = 110...but still...okay you're with me

OpenStudy (anonymous):

yup

OpenStudy (johnweldon1993):

So...now we have 3o - 2t = 110 and we now know that t = 1070 - o Since we know what 't' EQUALS....we can "substitute" this value into our equation for 't' 3o - 2t = 110 after replacing we have 3o - 2(1070 - o) = 110 Now we distribute that -2 into the parenthesis 3o - 2140 + 2o = 110 Still with me?

OpenStudy (johnweldon1993):

*I see you have posted another question...Since We were so close to finishing this one...I'll finish it up We were left with 3o - 2140 + 2o = 110 We want all of our numbers with 'o' on 1 side and every other number on the other side....so combine the 3o + 2o 5o - 2140 = 110 and add 2140 to both sides 5o - 2140 = 110 +2140 +2140 5o = 2250 and then we divide both sides by 5 because we want to solve for just 'o' right? not 5o.....so 5o = 2250 --------- 5 5 o = 450 And that would be your answer...I hope you understood this time